Yeetoofsawesome:

Two friends are playing tic-tac-toe. If Amy wins $\frac{3}{8}$ of the time, Lily wins $\frac{3}{10}$ of the time, and they tie the rest of the time, then what fraction of the time do they tie?

Yeetoofsawesome:

this one i need help

Mercury:

there are only three outcomes for tic-tac-toe: 1. Amy wins 2. Lily wins 3. They tie therefore, (fraction that Amy wins) + (fraction that Lily wins) + (fraction that they tie) = 1 therefore: fraction that they tie = 1 - (Amy's wins) - (Lily's wins) substituting the fractions from the problem: 1 - (3/8) - (3/10) = ?